# Advancing Brainwave Modeling with a Codebook-Based Foundation Model

### v1

@@ 1 Introduction
Brain-Computer Interface (BCI) technology enables a direct communication of the brain with the external world. This is attainable by analyzing brainwaves captured by electroencephalogram (EEG) recorders using advanced signal processing and, more recently, machine learning techniques. The roots of brainwave signal analysis lie in traditional neuroscience, where lengthy studies and hand-crafted features were once considered the gold standard [Bashashati et al. (2007), Handy (2009), Rao (2013), Nam et al. (2018), McFarland et al. (2006)]. However, as later research revealed, individual differences in brain activity (known as inter-subject variability) often prevent these features from generalizing effectively to real-world data, limiting their practicality in everyday BCI applications.
@@
As we have transitioned into the era of data-driven solutions, deep learning has reduced the need for manual feature extraction. Deep models can now automatically extract relevant features for various BCI paradigms, often achieving impressive performance [Lawhern et al. (2018); Santamaría-Vázquez et al. (2020); Song et al. (2023); Barmpas et al. (2023)]. Drawing inspiration from neuroscience Barmpas et al. (2024b), machine learning researchers have integrated domain insights into their models, addressing challenges such as inter-subject variability [Bakas et al. (2022); Wei et al. (2022)]. However, despite their successes, deep learning models typically require significant supervision and task-specific data collection, making the development process time consuming and resource-intensive.
@@
More recently, Large Foundation Models have transformed fields such as Computer Vision and Natural Language Processing with their remarkable results [Brown et al. (2020); Touvron et al. (2023); Mizrahi et al. (2023); Paraperas Papantoniou et al. (2024)]. These large models offer several advantages, including the ability to capture complex patterns through extensive self-supervised pre-training on diverse, unlabeled datasets. Therefore, they exhibit strong generalization capabilities, reducing the need for task-specific data collection and paradigm-specific model training. In the area of BCIs, Large Brainwave Foundation Models (LBMs) have emerged over the past few years [Jiang et al. (2024); Cui et al. (2024)]. These models have struggled to fully capture the rich information content of neural oscillations, a limitation that fundamentally constrains their performance and generalizability across diverse BCI tasks. Brain oscillations, the rhythmic electrical activity that forms the basis of EEG signals, are characterized by both amplitude and phase components. While amplitude reflects the power of neural activity, phase information encodes the precise timing of neural events and the synchronization between brain regions. Existing foundation models for EEG, including the recently introduced LaBraM Jiang et al. (2024), have made significant improvements by leveraging learning techniques from language and vision domains. These approaches typically apply vector-quantized Esser et al. (2020) neural spectrum prediction to convert continuous EEG signals into discrete neural codes. However, due to the way these models are trained, they suffer from a critical mathematical limitation: they fail to properly represent the circular topology of phase information.
@@ 2 Background
LaBraM Jiang et al. (2024) is a recently introduced unified EEG foundation model. Inspired by VQ-GAN Esser et al. (2020), LaBraM’s tokenizer is built on the vector-quantized neural spectrum prediction, trained to predict the Fourier spectrum of EEG signals. The encoding part is based on the neural transformer which receives patches of segmented EEG channels as input. Temporal convolution layers extract temporal features that are then combined with temporal and spatial embeddings and passed through a series of transformer layers Vaswani et al. (2023), with the incorporated modification as described in Dehghani et al. (2023):
$Attention(Q,V,K)=softmax(\frac{LN(Q)LN(K)^{T}}{\sqrt{d_{head}}})V$ (1)
where $d_{head}$ is the dimension of one head in multi-head attention, Q denotes the query, V denotes the value, K denotes the key and LN denotes Layer Norm Ba et al. (2016). The resulting patch representations are quantized into the neural codebook $\mathcal{V}$. These tokens are then passed into the decoding part (a shallow network) for reconstruction. The target objective is the reconstruction of the Fourier amplitude ($A$) and phase ($\phi$) of the EEG patches by minimizing the total loss:
@@
where $\hat{A}$ and $\hat{\phi}$ are the reconstructed amplitude and phase (direct amplitude $\mathcal{L}_{A}$ and direct phase $\mathcal{L}_{\phi}$ loss functions respectively) and $L_{Q}$ is the quantization loss as described in Esser et al. (2020).
While we agree with LaBraM’s core idea of encoding diverse neural activities into discrete neural tokens within a codebook $\mathcal{V}$ and using this codebook to train a large LBM, we identify several mathematical flaws in the codebook’s training process that we aim to address in this work.
@@ 3 Model
The proposed LaBraM++ model builds upon the original LaBraM introduced by Jiang et al. (2024). The primary contribution of this work is the development and training of an improved tokenizer with enhanced signal reconstruction capabilities. This improvement facilitates the training of a more effective LBM, resulting in superior performance across a range of tasks. Our key insight is that phase information (closely related to brain oscillations) should be represented in a way that respects the inherent circular structure of oscillatory dynamics. To achieve this, we adopt a sine/cosine phase encoding scheme that preserves the full informational content of neural oscillations while supporting stable gradient-based optimization.
3.1 Modified Tokenizer
@@
3.1.1 Description
Let $X\in\mathbb{R}^{C\times T}$ denote the input EEG signal, where $T$ is the number of time points and $C$ is the number of electrodes. In both the original LaBraM architecture and inherently in LaBraM++, the input EEG is first segmented into patches along the temporal dimension for each channel. Each patch is then passed through the neural transformer (as described in Section 2), which computes a patch representation $\mathbf{p}\in\mathbb{R}^{D}$, where $D$ is the embedding dimension.
As described previously, the tokenizer aims to learn a codebook of neural tokens, denoted as $\mathcal{V}$:
@@ 3.1.1 Description
Phase Information Issue: In the LaBraM architecture, the neural codebook $\mathcal{V}$ is trained via the reconstruction of the Fourier amplitude ($A$) and phase ($\phi$) of the EEG patches by minimizing the total loss $\mathcal{L}_{T}$, described in equation 2, which consists of the direct amplitude $\mathcal{L}_{A}$ and direct phase $\mathcal{L}_{\phi}$ loss function terms. However, $\mathcal{L}_{\phi}$ between actual $\phi$ and reconstructed $\hat{\phi}$ angles is not an ideal loss function due to the periodic nature of the Fourier phase. This can be seen most severely at the boundaries of the phase domain (±$\pm\pi$), where similar physiological states produce drastically different representations, causing severe information loss.
@@
$\mathcal{L}_{T}=\sum_{i}\underbrace{\left\|\hat{A_{i}}-A_{i}\right\|_{2}^{2}}_{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}\text{Amplitude loss - $\mathcal{L}_{A}$}}}+\underbrace{\left\|\sin{\hat{\phi_{i}}}-\sin{\phi_{i}}\right\|_{2}^{2}}_{{\color[rgb]{0.5,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0}\pgfsys@color@cmyk@stroke{0}{0}{1}{.5}\pgfsys@color@cmyk@fill{0}{0}{1}{.5}\text{Sine Phase loss - $\mathcal{L}_{sin}$}}}+\underbrace{\left\|\cos{\hat{\phi_{i}}}-\cos{\phi_{i}}\right\|_{2}^{2}}_{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\text{Cosine Phase loss - $\mathcal{L}_{cos}$}}}+\mathcal{L}_{Q}$ (5)
where $\hat{A}$, $\sin{\hat{\phi}}$ and $\cos{\hat{\phi}}$ are the reconstructed amplitude and phase sin/cos values respectively, and $L_{Q}$ is the quantization loss as described in Esser et al. (2020). The phases are essentially represented as unit vectors in 2D space, making the errors between them more geometrically meaningful.
@@ 3.2 Theoretical Basis
Limitations of Direct Phase Representation: The direct phase loss function (used in LaBraM) $\mathcal{L}_{\phi}=||\hat{\phi}-\phi||^{2}$ between actual $\phi$ and reconstructed $\hat{\phi}$ angles is discontinuous at the boundaries of the phase domain $[-\pi,\pi]$, creating optimization challenges. Consider the case where the predicted phase is $\hat{\phi}=\pi-\epsilon$ and the actual phase is $\phi=-\pi+\epsilon$, with $\epsilon>0$ small. Using the direct phase loss function:
$\mathcal{L}_{\phi}=||\hat{\phi}-\phi||^{2}=||2\pi-2\epsilon||^{2}=(2\pi-2\epsilon)^{2}$ (6)
As $\epsilon\rightarrow 0$, this approaches $4\pi^{2}$, despite representing angles that are close. Therefore, a small phase change, for example of just 2 degrees between $179^{\circ}$ and $-179^{\circ}$, can lead to a huge discontinuity shift. From a machine learning perspective, this problem creates fundamental optimization challenges. The discontinuity at phase boundaries leads to unstable gradients, poor convergence properties and ultimately suboptimal representation learning. This can make optimization unstable and lead to poor signal reconstruction.
Impact on Gradient-Based Optimization: Using the direct phase loss function $\mathcal{L}_{\phi}=||\hat{\phi}-\phi||^{2}$ between actual $\phi$ and reconstructed $\hat{\phi}$ angles leads to unstable gradients and poor optimization properties. For gradient-based optimization, we need to compute:
@@
As $\epsilon\to 0$, the magnitude of this gradient approaches $4\pi$, which is significantly larger than the gradients elsewhere in the phase domain. The large disparity in gradient magnitudes introduces several significant challenges during optimization. Abrupt changes in gradient values can result in numerical instability. In addition, gradient explosions can take place since excessively large gradients near the boundaries can cause disproportionately large optimization steps, potentially overshooting the optimal solution. In other words, these insights mean that neural networks trained with the direct phase loss function will struggle to learn accurate representations of phases near the ±$\pm\pi$ boundaries. This mathematical inconsistency creates a representation bottleneck that constrains model performance regardless of architecture size or training data volume.
@@
$\mathcal{L}_{r(\phi)}={\color[rgb]{0.5,0.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.5,0}\pgfsys@color@cmyk@stroke{0}{0}{1}{.5}\pgfsys@color@cmyk@fill{0}{0}{1}{.5}\mathcal{L}_{sin}}+{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathcal{L}_{cos}}=||\sin(\hat{\phi})-\sin(\phi)||^{2}+||\cos(\hat{\phi})-\cos(\phi)||^{2}=2-2\cos(\hat{\phi}-\phi)$ (11)
between actual $\phi$ and predicted angles $\hat{\phi}$ preserves more phase information than direct phase representation, particularly for phases near the ±$\pm\pi$ boundary.
@@
As $\epsilon\rightarrow 0$, this approaches 0. In the previous example, a phase change between $179^{\circ}$ and $-179^{\circ}$ is now translated to a small smooth difference since $cos(179^{\circ})\simeq cos(-179^{\circ})$.
@@ 4.1.1 Data Preprocessing
For a fair comparison, we aimed to pre-train our models with the same datasets used in the original LaBraM. While we have included the majority of the publicly available datasets, some could not be sourced, including the authors’ self-collected dataset. To compensate for this deficit, we also include $\sim$235 hours of our own self-collected motor data. The full breakdown of included and excluded data is given in Supplementary Material B.
@@ 4.1.3 Hyperparameters
LaBraM++ was trained using the same settings as the original LaBraM architecture (a detailed list of the training settings is provided in Supplementary Material A). The modified tokenizer was trained for 100 epochs with a neural codebook defined as $\mathcal{V}\in\mathbb{R}^{8192\times 64}$. Due to computational resource limitations, the base architecture of LaBraM was utilized for the core foundation model, which was trained for 50 epochs. Both models were trained on the datasets described in Supplementary Material B.
@@ 4.2 Comparison with LaBraM
The first step involves comparing LaBraM++ against the original LaBraM Jiang et al. (2024). As mentioned above, it is not possible to use the exact same datasets as in Jiang et al. (2024) for model pre-training. Therefore, for fair comparison we pre-trained both architectures LaBraM++ and LaBraM using the same datasets, pre-processing steps and implementation settings as described in sections 4.1.1, 4.1.2 and 4.1.3.
@@ 4.2.3 Fine-Tuning Performance
The next step is to compare the classification performances of these two models. We evaluated the performance of LaBraM++ against the original LaBraM architecture on full finetuning (training of the LBM and the added classification layer). Both models were evaluated in downstream classification tasks for the following five EEG datasets as described in the benchmark Lee et al. (2025): Motor paradigm in High Gamma Schirrmeister et al. (2017), a Working Memory dataset Pavlov et al. (2022), Physionet’s sleep staging dataset, Sleep-EDF Kemp et al. (2000) and Eyes Open vs Closed classification on the Physionet Motor dataset Schalk et al. (2004). These tasks were selected to capture a diverse range of BCI paradigms and the datasets were specifically chosen for their minimal spurious artifacts, reducing the likelihood of specious performance during training Lee et al. (2025).
Each model was trained for 20 epochs (to avoid overfitting) and evaluated using 10-fold subject-independent cross-validation, where samples were split on a subject level such that no subject would be present in both the training and validation sets. To perform the downstream tasks, untrained classification heads were added to the pre-trained LBMs before finetuning. The results of the comparisons are displayed in Table 1:
Table 1: Classification accuracy of LaBraM models after finetuning for 20 epochs ( denotes early stopping to avoid overfitting) with 10-fold cross validation. Bold values indicate best performance (per task or overall).
@@
Table 2: Classification accuracy of LaBraM++ variations after finetuning for 20 epochs ( denotes early stopping to avoid overfitting) with 10-fold cross validation. RA denotes Time-Embedding Right-Alignment and CAR denotes Common Average Reference. Bold and underlined values indicate best performance and next-best performance respectively (per task or overall).
@@ 4.3 Comparison with State-of-the-Art
Table 3: Classification accuracy of finetuned foundation models and standard deep learning architectures. Each trained/finetuned for 20 epochs ( denotes early stopping to avoid overfitting) with 10 fold cross-validation. Bold and underlined values indicate best performance and next-best performance respectively (per task or overall).
@@
We evaluated the performance of fine-tuned LaBraM++ in comparison to other LBMs, namely variations of NeuroGPT. By using the same benchmarking as described in section 4.1.3 and originally introduced in Lee et al. (2025). We perform the same full finetuning and use the same 10-fold subject-independent cross-validation for model evaluation. We chose the pre-trained NeuroGPT model and the encoder-only module of the pre-trained NeuroGPT model (as discussed in Cui et al. (2024) where the authors claim that fine-tuning the encoder alone produces similar or improved results over the full model).
As shown in Table 3, when full-finetuned, LaBraM++, despite being trained on less training data, it demonstrates competitive performance against other leading LBM models such as LaBraM and NeuroGPT. This suggests that if LaBraM++ was trained on the full list of datasets used by existing LBMs, it would likely yield even greater performance improvements, further reinforcing its potential as a strong and efficient brainwave foundation model.

### v2

@@ 1 Introduction
Brain-Computer Interface (BCI) technology enables a direct communication of the brain with the external world. This is attainable by analyzing brainwaves captured by electroencephalogram (EEG) recorders using advanced signal processing and, more recently, machine learning techniques. The roots of brainwave signal analysis lie in traditional neuroscience, where lengthy studies and hand-crafted features were once considered the gold standard [Bashashati et al., (2007), Handy, (2009), Rao, (2013), Nam et al., (2018), McFarland et al., (2006)]. However, as later research revealed, individual differences in brain activity (known as inter-subject variability) often prevent these features from generalizing effectively to real-world data, limiting their practicality in everyday BCI applications.
@@
As we have transitioned into the era of data-driven solutions, deep learning has reduced the need for manual feature extraction. Deep models can now automatically extract relevant features for various BCI paradigms, often achieving impressive performance [Lawhern et al., (2018); Santamaría-Vázquez et al., (2020); Song et al., (2023); Barmpas et al., (2023)]. Drawing inspiration from neuroscience Barmpas et al., 2024b , machine learning researchers have integrated domain insights into their models, addressing challenges such as inter-subject variability [Bakas et al., (2022); Wei et al., (2022)]. However, despite their successes, deep learning models typically require significant supervision and task-specific data collection, making the development process time consuming and resource-intensive.
@@
More recently, Large Foundation Models have transformed fields such as Computer Vision and Natural Language Processing with their remarkable results [Brown et al., (2020); Touvron et al., (2023); Mizrahi et al., (2023); Paraperas Papantoniou et al., (2024)]. These large models offer several advantages, including the ability to capture complex patterns through extensive self-supervised pre-training on diverse, unlabeled datasets. Therefore, they exhibit strong generalization capabilities, reducing the need for task-specific data collection and paradigm-specific model training. In the area of BCIs, Large Brainwave Foundation Models (LBMs) have emerged over the past few years [Jiang et al., (2024); Cui et al., (2024)]. These models have struggled to fully capture the rich information content of neural oscillations, a limitation that fundamentally constrains their performance and generalizability across diverse BCI tasks. Brain oscillations, the rhythmic electrical activity that forms the basis of EEG signals, are characterized by both amplitude and phase components. While amplitude reflects the power of neural activity, phase information encodes the precise timing of neural events and the synchronization between brain regions. Existing foundation models for EEG, including the recently introduced LaBraM Jiang et al., (2024), have made significant improvements by leveraging learning techniques from language and vision domains. These approaches typically apply vector-quantized Esser et al., (2020) neural spectrum prediction to convert continuous EEG signals into discrete neural codes. However, due to the way these models are trained, they suffer from a critical mathematical limitation: they fail to properly represent the circular topology of phase information.
@@ 2 Background
LaBraM Jiang et al., (2024) is a recently introduced unified EEG foundation model. Inspired by VQ-GAN Esser et al., (2020), LaBraM’s tokenizer is built on the vector-quantized neural spectrum prediction, trained to predict the Fourier spectrum of EEG signals. The encoding part is based on the neural transformer which receives patches of segmented EEG channels as input. Temporal convolution layers extract temporal features that are then combined with temporal and spatial embeddings and passed through a series of transformer layers Vaswani et al., (2023), with the incorporated modification as described in Dehghani et al., (2023). The resulting patch representations are quantized into the neural codebook $\mathcal{V}$. These tokens are then passed into the decoding part (a shallow network) for reconstruction. The target objective is the reconstruction of the Fourier amplitude ($A$) and phase ($\phi$) of the EEG patches by minimizing the total loss:
@@
where $\hat{A}$ and $\hat{\phi}$ are the reconstructed amplitude and phase (direct amplitude $\mathcal{L}_{A}$ and direct phase $\mathcal{L}_{\phi}$ loss functions respectively) and $L_{Q}$ is the quantization loss as described in Esser et al., (2020). While we agree with LaBraM’s core idea of encoding diverse neural activities into discrete neural tokens within a codebook $\mathcal{V}$ and using this codebook to train a large LBM, we identify several mathematical flaws in the codebook’s training process that we aim to address in this work.
@@ 3 Model
The proposed LaBraM++ model builds upon the original LaBraM introduced by Jiang et al., (2024). The primary contribution of this work is the development and training of an improved tokenizer with enhanced signal reconstruction capabilities. This improvement facilitates the training of a more effective LBM, resulting in superior performance across a range of tasks. Our key insight is that phase information (closely related to brain oscillations) should be represented in a way that respects the inherent circular structure of oscillatory dynamics. To achieve this, we adopt a sine/cosine phase encoding scheme that preserves the full informational content of neural oscillations while supporting stable gradient-based optimization.
@@
3.1 Description
Let $X\in\mathbb{R}^{C\times T}$ denote the input EEG signal, where $T$ is the number of time points and $C$ is the number of electrodes. In both the original LaBraM architecture and inherently in LaBraM++, the input EEG is first segmented into patches along the temporal dimension for each channel. Each patch is then passed through the neural transformer (as described in Section 2), which computes a patch representation $\mathbf{p}\in\mathbb{R}^{D}$, where $D$ is the embedding dimension. As described previously, the tokenizer aims to learn a codebook of neural tokens, denoted as $\mathcal{V}$:
@@ 3.1 Description
where $l_{2}$ represents normalization.
Phase Information Issue: In the LaBraM architecture, the neural codebook $\mathcal{V}$ is trained via the reconstruction of the Fourier amplitude ($A$) and phase ($\phi$) of the EEG patches by minimizing the total loss $\mathcal{L}_{T}$, described in equation 1, which consists of the direct amplitude $\mathcal{L}_{A}$ and direct phase $\mathcal{L}_{\phi}$ loss function terms. However, $\mathcal{L}_{\phi}$ between actual $\phi$ and reconstructed $\hat{\phi}$ angles is not an ideal loss function due to the periodic nature of the Fourier phase. This can be seen most severely at the boundaries of the phase domain ($\pm\pi$), where similar physiological states produce drastically different representations, causing severe information loss (Section 3.2).
@@
[…]
where $\hat{A}$, $\sin{\hat{\phi}}$ and $\cos{\hat{\phi}}$ are the reconstructed amplitude and phase sin/cos values respectively, and $L_{Q}$ is the quantization loss as described in Esser et al., (2020). The phases are essentially represented as unit vectors in 2D space, making the errors between them more geometrically meaningful (Section 3.2).
@@ 3.2 Theoretical Basis
[…]
As $\epsilon\rightarrow 0$, this approaches $4\pi^{2}$, despite representing angles that are close. Therefore, a small phase change, for example of just 2 degrees between $179^{\circ}$ and $179^{\circ}$, can lead to a huge discontinuity shift. From a machine learning perspective, this problem creates fundamental optimization challenges. The discontinuity at phase boundaries leads to unstable gradients, poor convergence properties and ultimately suboptimal representation learning. This can make optimization unstable and lead to poor signal reconstruction.
[…]
@@
As $\epsilon\to 0$, the magnitude of this gradient approaches $4\pi$, which is significantly larger than the gradients elsewhere in the phase domain. The large disparity in gradient magnitudes introduces several significant challenges during optimization. Abrupt changes in gradient values can result in numerical instability. In addition, gradient explosions can take place since excessively large gradients near the boundaries can cause disproportionately large optimization steps, potentially overshooting the optimal solution. In other words, these insights mean that neural networks trained with the direct phase loss function will struggle to learn accurate representations of phases near the $\pm\pi$ boundaries. This mathematical inconsistency creates a representation bottleneck that constrains model performance regardless of architecture size or training data volume.
@@
[…]
between actual $\phi$ and predicted angles $\hat{\phi}$ preserves more phase information than direct phase representation, particularly for phases near the $\pm\pi$ boundary.
@@
As $\epsilon\rightarrow 0$, this approaches 0. In the previous example, a phase change between $179^{\circ}$ and $179^{\circ}$ is now translated to a small smooth difference since $cos(179^{\circ})\simeq cos(-179^{\circ})$.
@@ 4.1.1 Data Preprocessing
For a fair comparison, we aimed to pre-train our models with the same datasets used in the original LaBraM. While we have included the majority of the publicly available datasets, some could not be sourced, including the authors’ self-collected dataset. To compensate for this deficit, we also include $\sim$235 hours of our own self-collected motor data ***EEG Data Collection taken place at Imperial College London : https://mybraincommands.doc.ic.ac.uk. The full breakdown of included and excluded data is given in Appendix B.
@@ 4.1.3 Hyperparameters
LaBraM++ was trained using the same settings as the original LaBraM architecture (a detailed list of the training settings is provided in Appendix A). The modified tokenizer was trained for 100 epochs with a neural codebook defined as $\mathcal{V}\in\mathbb{R}^{8192\times 64}$. Due to computational resource limitations, the base architecture of LaBraM was utilized for the core foundation model, which was trained for 50 epochs. Both models were trained on the datasets described in Appendix B.
@@ 4.2 Comparison with LaBraM
The first step involves comparing LaBraM++ against the original LaBraM Jiang et al., (2024). Since a self-collected dataset has been used, it is not possible to use the exact same datasets as in Jiang et al., (2024) for model pre-training. Therefore, for fair comparison we pre-trained both architectures LaBraM++ and LaBraM using the same datasets, pre-processing steps and implementation settings as described in sections 4.1.1, 4.1.2 and 4.1.3.
@@ 4.2.3 Fine-Tuning Performance
The next step is to compare the classification performances of these two models. We evaluated the performance of LaBraM++ against the original LaBraM architecture on full finetuning (training of the LBM and the added classification layer). Both models were evaluated in downstream classification tasks for the following four EEG datasets as described in the benchmark Lee et al., (2025). These tasks were selected to capture a diverse range of BCI paradigms and the datasets were specifically chosen for their minimal spurious artifacts, reducing the likelihood of spurious performance during training Lee et al., (2025). Each model was trained for 20 epochs (to avoid overfitting) and evaluated using 10-fold subject-independent cross-validation, where samples were split on a subject level such that no subject would be present in both the training and validation sets. To perform the downstream tasks, untrained classification heads were added to the pre-trained LBMs before finetuning. The results of the comparisons are displayed in Table 1.
Table 1: Classification balanced accuracy of LaBraM and LaBraM++ models after finetuning for 20 epochs ( denotes early stopping to avoid overfitting) with 10-fold cross validation. Bold values indicate best performance (per task or overall).
@@
Table 2: Classification balanced accuracy of LaBraM++ variations after finetuning for 20 epochs ( denotes early stopping to avoid overfitting) with 10-fold cross validation. RA denotes Time-Embedding Right-Alignment and CAR denotes Common Average Reference. Bold and underlined values indicate best performance and next-best performance respectively (per task or overall).
@@ 4.3 Comparison with State-of-the-Art
We evaluated the performance of fine-tuned LaBraM++ in comparison to other fine-tuned LBMs, namely variations of NeuroGPT Cui et al., (2024), BIOT Yang et al., (2023), EEGPT Wang et al., (2024) and CBraMod Wang et al., (2025). By using the same benchmarking as described in section 4.2.3 and originally introduced in Lee et al., (2025), we perform the same full finetuning and use the same 10-fold subject-independent cross-validation for the evaluation of all models. As shown in Table 3, when full-finetuned, LaBraM++ demonstrates competitive performance against other leading finetuned LBM models. This suggests that training LaBraM++ on a broader range of datasets used by existing LBMs could further enhance its performance, reinforcing its potential as a strong and efficient brainwave foundation model.
Table 3: Classification balanced accuracy of finetuned foundation models. Each trained/finetuned for 20 epochs ( denotes early stopping to avoid overfitting) with 10 fold cross-validation. Bold and underlined values indicate best performance and next-best performance respectively (per task or overall). RA denotes Time-Embedding Right-Alignment and CAR denotes Common Average Reference.